two masses 5kg and 2kg connected by a rope .The 5kg mass experiences a kinetic friction of 9.8N between it and the table. Determine the i) acceleration of the system ii) tension in the rope


1
Expert's answer
2021-09-07T09:50:31-0400

Given:

"m_1=5\:\rm kg"

"m_2=2\:\rm kg"

"F_f=9.8\:\rm N"


The Newton's second law says

"m_1a=T-F_f\\
m_2a=m_2g-T"

So,

(i)

"(m_1+m_2)a=m_2g-F_f\\
a=\frac{m_2g-F_f}{m_1+m_2}"

"a=\frac{2*9.8-9.8}{5+2}=1.4\:\rm m/s^2"

(ii)

"T=m_2g-m_2a=2*(9.8-1.4)=16.8\:\rm N"